OpenStudy (shyannep913):

help what is the equation of the line x-3y=18 in slope-intercept form?

HanAkoSolo (jamierox4ev3r):

First of all, do you know what slope-intercept form is?

OpenStudy (shyannep913):

yes

HanAkoSolo (jamierox4ev3r):

so then, what variable are we solving for here?

OpenStudy (shyannep913):

y

HanAkoSolo (jamierox4ev3r):

correct. So if we're given x-3y=18, what is our first step in solving?

OpenStudy (shyannep913):

subtract x from each side?

HanAkoSolo (jamierox4ev3r):

yes. So what would your equation look like after doing that?

OpenStudy (shyannep913):

3y=x-18?

HanAkoSolo (jamierox4ev3r):

ah, but don't forget not to mess up your signs.

HanAkoSolo (jamierox4ev3r):

Here's how the subtraction should look like: x-3y=18 -x -x -------- -3y= -x+18 ^ I know it's not pretty, but we have to keep our signs. Does this make sense?

OpenStudy (shyannep913):

yes

HanAkoSolo (jamierox4ev3r):

okay, so what is our final step?

OpenStudy (shyannep913):

3y/3=-x-18/3?

HanAkoSolo (jamierox4ev3r):

Ah, like I said, don't forget to keep signs constant. But you are correct in dividing both sides by 3

HanAkoSolo (jamierox4ev3r):

it would look something more like this: -3y= -x+18 /-3 /-3 /-3 ---------- y= ?

HanAkoSolo (jamierox4ev3r):

can you finish the rest? :)

OpenStudy (shyannep913):

yes

OpenStudy (shyannep913):

thank you

HanAkoSolo (jamierox4ev3r):

no problem